I am working on this question and i cant seem to get please can i get help on this question ..
Which of the following points would fall on the line produced by this point-slope form equation y + 5 = 2(x – 1) when graphed?
a.(0,-7)
b.(2,-3)
c.(0,-3)
d.(-7,2)-i picked

Answer by stanbon(75887)

Which of the following points would fall on the line produced by this point-slope form equation y + 5 = 2(x – 1) when graphed?
a.(0,-7)
-7 + 5 = 2(0-1)
-2 = 2*-1 = -2
---------------------
b.(2,-3)
-3 + 5 = 2(2-1)
2 = 2(1)
2 = 2
--------------------
c.(0,-3)
-3+5 = 2(0-1)
2 = -2 ----------No
--------------------
d.(-7,2)-i picked
2+5 = 2(-7-1)
7 = 2(-8)
7 = -16--------No
===================
Cheers,
Stan H.
===================